Can you do number 3
Construct and calculate a theoretical titration cu
Answer #1

Answer (3)

  • First of all 6.881 g of MgCl2.6H2O and NaCl dissolved in 500 mL. One tenth of that (50 mL) gave 0.5923 g of AgCl (mol. wt. = 143.4 g) so 0.5923/143.4 = 0.004132 mole AgCl and the same number of moles of Cl.
  • Mg2P2O7 has mol. wt. 222.6 g so 0.1796/222.6 = 0.0008068 moles of Mg2P2O7 which means 0.001614 mole Mg (two mole Mg per mol Mg2P2O7)
  • So 500 mL contained 10x that moles Mg = 0.01614 which means that solid contained 0.01614 moles MgCl2.6H2O (mol. wt. = 203.3) so 0.01614*203.3 = 3.281 g which leaves 3.600 g of NaCl.
  • 3.600 g NaCl are (3.600/58.5) moles = 0.06160 moles of NaCl which means the sample contained 0.0616 +( 2 x 0.01614) moles Cl = 0.09387 moles which is not the answer for the moles of Cl.
  • If the former, moles Cl in NaCl = 0.4132 - 2*0.01614 = 0.009044 moles so g NaCl = 0.009044*58.5 = 0.5287 g.
  • That suggests 3.281 g are magnesium chloride; 0.5287 are NaCl. %ages are 47.68%and 7.68%